# LOGIN: A Large Language Model Consulted Graph Neural Network Training Framework

### v1

@@ 4.2. Performance Comparison (RQ1)
Table 2. Performance Comparison for node classification on homophilic and heterophilic graphs. The best results and the second best results among the classic GNNs with and without LOGIN and the advanced models are bold and underlined respectively.
@@
To answer the first research question, we evaluate our proposed L O G IN framework on three homophilic datasets and three heterophilic datasets. The prediction accuracy scores and standard errors are reported in Table 2. Comparing fundamental baselines within L O G IN to the advanced GNNs with complex designs, we have the following observations.
@@
Firstly, our method consistently outperforms the vanilla baselines trained with the original node features or the LM-finetuned node embeddings in most cases. There are only two exceptions with the Cornell dataset. Since Cornell is a relatively small dataset with only 39 nodes in the test set, as indicated by the standard errors, the experimental randomness is quite high with this small dataset. Nevertheless, L O G IN still helps improve the performance of MixHop on Cornell by 4.6% with a low deviation. Apart from this exception, all listed fundamental baselines trained within our L O G IN framework exceed the vanilla ones in node classification accuracy on the other five datasets, demonstrating the effectiveness of integrating LLMs as consultants into the GNN training process.
@@
Secondly, we are able to attain performance comparable to that of advanced GNNs by training fundamental models within the L O G IN framework. It is noteworthy that on PubMed and Texas, respectively known as benchmarks for homophilic and heterophilic graphs, we achieve the highest prediction accuracy among all the compared methods. This finding verifies the generalizability of our method on graphs with distinctive characteristics. For Cora, Arxiv-23, Wisconsin, and Cornell, L O G IN achieve remarkable performance on par with the intricately designed GNNs as well.
@@
Thirdly, it draws our attention that regardless of the feature type and training paradigm we apply, MLP reveals great potential in the node classification task on heterophilic graphs. We believe that the TAGs with considerably high heterophily may be better regarded as natural language data rather than graph-structured data, for the links contained in these datasets seem to only carry noticeably limited information. Additionally, MLP also demonstrates remarkable ability on homophilic graphs, inspiring further research on how to leverage simple existing tools to achieve superb capability.
@@ 4.3. Ablation Study (RQ2)
Regarding the structure refinement component, L O G IN without edge pruning exhibits a significant performance decrease compared to the models with the complete coping mechanism, as shown in Fig. 3 (b). This observation aligns with our intuition, as in heterophilic graphs, the principal challenge for conventional neural networks in achieving generalization stems from the distinctiveness of their structural characteristics.
@@ 4.4. Case Study (RQ3)
We present node 356 from Cora as a representative example, whose ground truth label is Neural Networks. Unlike other papers, the title and abstract of this paper do not feature its label as a term explicitly, which also poses challenges for human classification. Besides, node 356 only has two one-hop neighbors, one of which is labeled differently as Probabilistic Method. This discrepancy may lead to failure in GNNs when processing this node. Nevertheless, thanks to the comprehensive understanding of the paper content facilitated by the parametric knowledge of LLMs, accurate predictions and concise rationales are generated. Consequently, the subsequent semantic enhancement of node features significantly contributes to the final prediction of the GNN. The consultation process is elaborated in Fig. 4.
Figure 4. Case study of Cora 356.
@@
Node 62 from Wisconsin represents a web page of a course, with content that is clear enough for humans to identify as a course homepage. However, due to its misleading neighborhood, where all nodes except itself in its 2-hop ego-graph do not represent course, among which 3 out of 4 are web pages of students, pre-trained GNNs cannot directly classify node 8 accurately. Moreover, the LLM consultant also provides the incorrect classification, along with an illogical rationale mentioning GNN prediction as in Fig. 5. This leads to the pruning of all the links of node 62, which contributes to structure denoising that aids the re-trained GNN in making the right choice.
Figure 5. Case study of Wisconsin 62.
The studies of two interesting cases we encountered in experiments highlight the solidity and efficacy of L O G IN when handling various scenarios, which are consistent with our motivation of designing an LLM-fault-utilized strategy, turning dust into gold.
@@ 4.5. Extensive Study (RQ4)
To answer the fourth question, we work on extended studies with more advanced LLMs, namely vicuna-v1.5-13b (Zheng et al., 2023) and GPT 3.5-turbo-0125 (Brown et al., 2020). Due to constraints related to computational resources and OpenAI API calling, we provide results solely for Cora, based on two runs as presented in Table 3.
@@ 4.6. Comparison among LLM-based Paradigms (RQ5)
Last but not least, to answer the fifth research question, we investigate the LLMs-as-Predictors, LLMs-as-Enhancers, and our LLM-as-Consultants paradigms by testing typical methods derived from each. For LLMs-as-Predictors, we prompt vicuna-v1.5-7b (Zheng et al., 2023) to directly get predictions. For LLMs-as-Enhancers, we adopt the TAPE method (He et al., 2023) equipped with llama2-13b-chat(Touvron et al., 2023).
@@
5. Related Works
5.1. GNN Variants
Early proposed GNNs(Kipf and Welling, 2016; Hamilton et al., 2017a; Velickovic et al., 2017) are primarily designed for the most typical graph-structured data, such as citation networks, which are characterized as homophilic, homogeneous, and class-balanced graphs. Consequently, when generalized to heterogeneous (Zhang et al., 2019; Wang et al., 2019a), heterophilic (Bo et al., 2021; Abu-El-Haija et al., 2019), and class-imbalanced graphs (Wang et al., 2019b; Liu et al., 2021a), conventional GNNs suffer from severe performance degradation.
To resolve such present challenges, researchers have been driven to tailor conventional GNNs to specific scenarios. For example, in the case of heterophily, the designs are generally categorized into neighbor extension and GNN architecture refinement (Zheng et al., 2022).
For neighbor expansion, there are two subcategories, namely high-order neighbor mixing and potential neighbor discovery. High-order neighbor mixing, exemplified by MixHop and H2GCN (Zhu et al., 2020), enables GNNs to aggregate information across multi-hop ego-graphs. UGCN (Jin et al., 2021) also utilizes the two-hop neighbor set and further confines the set scale. Geom-GCN (Pei et al., 2020) discovers potential neighbors by discerning neighbor connections based on geometric relationships. Moreover, integrating class-aware information into the propagation process enables GNNs to focus more on intra-class nodes, as demonstrated in NL-GNN (Liu et al., 2021b), GPNN (Wang et al., 2022).
The GNN architecture refinement approach is divided into three types: adaptive message aggregation, ego-neighbor separation, and inter-layer combination. Regarding message aggregation, FAGCN (Bo et al., 2021) and ACMGCN (Luan et al., 2022) design adaptive filters for spectral GNNs to capture high-frequency signals. For spatial GNNs, the aggregation strategies are designed based on graph structure characteristics, as in DMP (Yang et al., 2021) and CPGNN (Zhu et al., 2021). Ego-neighbor separation helps GNNs produce distinguishable node embeddings. H2GCN (Zhu et al., 2020) exemplifies this method by excluding self-loop connections for non-mixing node embedding learning. Inter-layer combination methods enhance GNNs by conducting layer-wise operations. JKNet (Xu et al., 2018) is the first to propose the idea of combining intermediate representations from each layer, allowing flexible adaptation of neighborhood ranges to individual nodes. H2GCN (Zhu et al., 2020) and GCNII (Chen et al., 2020) further utilize all intermediate representations and only the first layer’s node embedding at each layer with the initial residual connection to realize inter-layer combination, respectively.
5.2. LLMs for Graphs
Regarding the former paradigm, NLGraph (Wang et al., 2023a), GPT4Graph (Guo et al., 2023), GraphQA (Fatemi et al., 2023) and GraphText (Zhao et al., 2023b) attempt to harness the zero-shot and few-shot ability, along with the in-context learning ability to solve graph tasks by describing the graph structural topology in natural language. However, the global graph structure is too complex to be compressed in a token-limited prompt, thereby utilizing LLMs solely only achieves the performance far from desired. Besides directly prompting LLMs, GraphLLM (Chai et al., 2023) conducts prefix tuning on an open-source LLM by concatenating graph-specific prefixes to the attention layers of a pre-trained LLM. InstructGLM (Ye et al., 2023) and GraphGPT (Tang et al., 2023) adapt LLMs for graph downstream tasks through instruction tuning, employing natural language and a graph-text aligner to express graph structural information, respectively. Nevertheless, neglecting the authenticated power of existing GNNs results in only moderate performance or substantial computing resource consumption.
Regarding the latter paradigm, TAPE (He et al., 2023) and Graph-LLM (Chen et al., 2023) prompt LLMs to interpret original texts attached to the nodes, and provide GNNs with semantically enhanced node features. LLM-TSE (Sun et al., 2023) explicitly instructs LLMs to produce a similarity score for two texts of two nodes, which subsequently leads to edge pruning. This methodology employs LLMs solely as a one-time data preprocessor before GNN training.
In addition, there are also works (Liu et al., 2023c; Huang and Chang, 2022; Sun et al., 2023; Liu et al., 2023a) in the field of graph prompting that apply the principles of LLMs to the domain of graph machine learning. Inspired by the route of how LLMs unify various natural language tasks, GraphPrompt (Liu et al., 2023c), Prodigy (Huang and Chang, 2022), All-in-one (Sun et al., 2023), and OFA (Liu et al., 2023a) aim to develop approaches for graph-structured prompt learning analogous to natural language prompt learning, with the objective of further unifying all downstream tasks in graph learning.
6. Conclusion and Future Work
In this work, we propose a new paradigm of leveraging LLMs for graph tasks, coined “LLMs-as-Consultants”. Following this paradigm, our L O G IN framework empowers interactive LLM consultation in the GNN training process. We identify crucial nodes in the GNN pre-training stage, and prompt LLMs with rich semantic and topological information compression, parse LLMs’ responses in a not only fault-tolerant but also fault-utilized way to enhance GNN re-training. Extensive experiments on both homophilic and heterophilic graphs illustrate the validity and versatility of our proposed LLMs-as-Consultants paradigm.
Appendix A Example Prompts
We display 2 example node prompts on datasets Cora and Wisconsin as in Table 5 for reproducibility.
Datasets Prompts
Cora Given a citation graph: { ”id”: ”cora_356”, ”graph”: { ”node_idx”: 356, ”node_list”: [356, 190, 510, 519, 2223, 192], ”one_hop_neighbors”: [190, 510], ”two_hops_neighbors”: [192, 519, 2223], ”node_label”: [”Neural_Networks”, ”Neural_Networks”, ”Probabilistic_Methods”, ”Neural_Networks”, ”Neural_Networks”, ”Neural_Networks”] }}, where the 0th node is the target paper, with the following information: Title: [Title Text], Abstract: [Abstract Text]. And in the ’node label’ list / ’GNN-predicted node label’ list, you’ll find the human-annotated / GNN-predicted subcategories corresponding to the neighbors within two hops of the target paper as per the ’node_list’.
Wisconsin Given a webpage link graph: { ”id”: ”wisconsin_62”, ”graph”: { ”node_idx”: 62, ”node_list”: [62, 166, 189, 165, 84], ”one_hop_neighbors”: [166, 189], ”two_hops_neighbors”: [84, 165], ”node_label”: [”course”, ”student”, ”student”, ”student”, ”faculty”] }}, where the 0th node is the target webpage, with the following content: [webpage content text]. And in the ’node label’ list / ’GNN-predicted node label’ list, you’ll find the human-annotated / GNN-predicted subcategories corresponding to the neighbors within two hops of the target paper as per the ’node_list’.
Question: Which category does this webpage belong to? Give the most likely category of this webpage directly, choosing from ”course”, ”faculty”, ”student”, ”project”, ”staff”. Ensure that your response can be parsed by Python json, using the following format as an example: {”classification result”: ”student”, ”explanation”: ”your explanation for your classification here”,} Ensure that the classification result must match one of the given choices.
Table 5. Example Prompts on Cora and Wisconsin
Appendix B more nodes to consult LLM
@@ Appendix A Example Prompts
MLP+LOGIN+extra 10% nodes 0.7147 ± 0.0318 0.8605 ± 0.0493
SAGE+LOGIN 0.8727 ± 0.0219 0.7434 ± 0.0930
SAGE+LOGIN+extra 10% nodes 0.8789 ± 0.0310 0.7527 ± 0.0698

### v2

@@ 6.2. Performance Comparison (RQ1)
Table 2. Performance comparison for node classification on homophilic and heterophilic graphs. The best results and the second best results among the classic GNNs with and without L
O
G
IN
and the advanced models are bold and underlined respectively.
@@
To answer the first research question, we evaluate our proposed L O G IN framework on three homophilic datasets and three heterophilic datasets. The prediction accuracy scores and standard errors are reported in Table 2. We have the following observations.
@@
Firstly, our method consistently outperforms the vanilla baselines trained with the original node features or the LM-finetuned node embeddings in most cases. There are only two exceptions in the Cornell dataset. Since Cornell is a relatively small dataset with only 39 nodes in the test set, as indicated by the standard errors, the experimental randomness is quite high with this small dataset. Nevertheless, L O G IN still helps improve the performance of MixHop on Cornell by 4.6% with a low deviation. Apart from this exception, all listed fundamental baselines trained within our L O G IN framework exceed the vanilla ones in node classification accuracy on the other five datasets, demonstrating the effectiveness of integrating LLMs as consultants into the GNN training process.
@@
Secondly, we are able to attain performance comparable to that of advanced GNNs by training fundamental models within the L O G IN framework. It is noteworthy that on PubMed and Texas, respectively known as benchmarks for homophilic and heterophilic graphs, we achieve the highest prediction accuracy among all the compared methods. This finding verifies the generalizability of our method on graphs with distinctive characteristics. For Cora, Arxiv-23, Wisconsin, and Cornell, L O G IN achieve remarkable performance on par with the SOTA GNNs such as H2GCN and SSP as well.
@@
Thirdly, it draws our attention that regardless of the feature type and training paradigm we apply, MLP reveals great potential in the node classification task on heterophilic graphs. We believe that the TAGs with considerably high heterophily may be better regarded as natural language data rather than graph-structured data, for the links contained in these datasets seem to only carry noticeably limited information. Additionally, our L O G IN can clearly enhance MLP, inspiring further research on how to leverage simple existing tools to achieve superb capability with the help of the LLM-as-Consultants paradigm.
@@ 6.3. Ablation Study (RQ2)
Regarding the structure refinement component, L O G IN without edge pruning exhibits a significant performance decrease compared to the models with the complete coping mechanism, as shown in Fig. 3 (b). This observation aligns with our intuition, as in heterophilic graphs, the principal challenge for conventional GNNs in achieving generalization stems from the distinctiveness of their structural characteristics.
@@ 6.4. Case Study (RQ3)
We present node 356 from Cora as a representative example, whose ground truth label is Neural Networks. Unlike other papers, the title and abstract of this paper do not feature its label as a term explicitly, which also poses challenges for human classification. Besides, node 356 only has two one-hop neighbors, one of which is labeled differently as Probabilistic Method. This discrepancy may lead to failure in GNNs when processing this node. Nevertheless, thanks to the comprehensive understanding of the paper content facilitated by the parametric knowledge of LLMs, accurate prediction and a concise rationale are generated. Consequently, the subsequent semantic enhancement of node features significantly contributes to the final prediction of the GNN. The consultation process is elaborated in Fig. 4.
Figure 4. Case study of node 356 in Cora. Node 356 was first misclassified by GNN and selected as an uncertain node. The LLM consultation classified node 356 right along with a reasonable explanation. Then, the node representation was updated with the explanation. Finally, the retrained GNN managed to classify node 356 correctly.
@@
Node 62 from Wisconsin represents a web page of a course, with content that is clear enough for humans to identify as a course homepage. However, due to its misleading neighborhood, where all nodes except itself in its 2-hop ego-graph do not represent course, among which 3 out of 4 are web pages of students, pre-trained GNNs cannot directly classify node 8 accurately. Moreover, the LLM consultant also provides the incorrect classification, along with an illogical rationale mentioning GNN prediction as in Fig. 5. This leads to edge pruning around node 62, which contributes to structure denoising that aids the re-trained GNN in making the right choice.
Figure 5. Case study of node 62 in Wisconsin. Node 62 was first misclassified by GNN and selected as an uncertain node. The LLM consultation classified node 62 wrong along with a self-contradictory explanation, which L
O
G
IN
left out. Then, edge pruning was performed around node 62. Finally, the retrained GNN managed to classify node 62 correctly.
The studies of two interesting cases we encountered in experiments highlight the solidity and efficacy of L O G IN when handling various scenarios, which are consistent with our motivation of designing an LLM-fault-utilized strategy for feedback.
@@ 6.5. Extensive Study (RQ4)
To answer the fourth question, we work on extended studies with more advanced LLMs, namely vicuna-v1.5-13b (Zheng et al., 2023) and GPT 3.5-turbo-0125 (Brown et al., 2020). Due to the constraints related to computational resources and OpenAI API calling, we provide results solely for Cora, based on two runs as presented in Table 3.
@@ 6.6. Comparison among LLM-based Paradigms (RQ5)
To answer the fifth research question, we investigate the LLMs-as-Predictors, LLMs-as-Enhancers, and our LLM-as-Consultants paradigms by testing typical methods derived from each. For LLMs-as-Predictors, we prompt vicuna-v1.5-7b (Zheng et al., 2023) to directly get predictions. For LLMs-as-Enhancers, we adopt the TAPE method (He et al., 2023) equipped with llama2-13b-chat(Touvron et al., 2023).
@@
6.7. Consulting with More Nodes (RQ6)
To address the sixth research question, we conducted experiments on the Cora and Wisconsin datasets using MLP and GraphSAGE to investigate the impact of consulting LLMs with more nodes. Our findings demonstrate that incorporating more nodes in the consultation process with LLMs can lead to significant performance improvements. The results are in Table 5.
Through these experiments, we increased the number of nodes involved in the LLM consultation by 10% each. The results on two example homophilic and heterophilic graphs show that as the number of consulted nodes increases, there are corresponding improvements in the overall performance of the model. This suggests that LLMs, when provided with a broader context through more extensive node consultation, can offer more insightful responses, thereby refining the model’s predictions and decisions. These findings underscore the potential benefits of leveraging LLMs more comprehensively, highlighting the effectiveness of the node consultation process to fully exploit the capabilities of LLMs in improving GNN performance.
@@ 6.7. Consulting with More Nodes (RQ6)
MLP+LOGIN+10% nodes 0.7147 ± 0.0318 0.8605 ± 0.0493
GrraphSAGE+LOGIN 0.8727 ± 0.0219 0.7434 ± 0.0930
GraphSAGE+LOGIN+10% nodes 0.8789 ± 0.0310 0.7527 ± 0.0698
6.8. L O G IN with Advanced GNNs (RQ7)
To address the seventh research question, we performed experiments incorporating APPNP and GCNII with L O G IN to demonstrate that L O G IN can also enhance the performance of advanced GNNs.
The results in Table 6 indicate notable performance gains on both Cora and Wisconsin. These enhancements underscore the versatility and broad applicability of L O G IN in enhancing advanced GNNs as well, reinforcing its potential as a general solution for boosting various types of GNNs. These experiments provide evidence that L O G IN is generally effective and not limited to improving specific types of GNNs.
Table 6. Performance comparison of APPNP and GCNII trained with or without the L
O
G
IN
framework.
Method Cora Wisconsin
APPNP 0.8539 ± 0.0477 0.7969 ± 0.0143
APPNP+LOGIN 0.8614 ± 0.0332 0.7973 ± 0.0202
GCNII 0.8833 ± 0.0027 0.7847 ± 0.0068
GCNII+LOGIN 0.8879 ± 0.0105 0.7903 ± 0.0121
7. Conclusion and Future Work
In this work, we propose a new paradigm of leveraging LLMs for graph tasks, coined “LLMs-as-Consultants”. Following this paradigm, our L O G IN framework empowers interactive LLM consultation in the GNN training process. We identify uncertain nodes in the GNN pre-training stage, prompt LLMs with rich semantic and topological information compression, and parse LLMs’ responses in a not only fault-tolerant but also fault-utilized way to enhance GNN re-training. Extensive experiments on both homophilic and heterophilic graphs illustrate the validity and versatility of our proposed LLMs-as-Consultants paradigm.
Moreover, a significant challenge that remains is determining how to efficiently involve large-scale nodes in the LLM consultation process. This is crucial for improving the scalability and practicality of the LLMs-as-Consultants paradigm. Addressing this issue involves developing methods that can handle the extensive computational requirements and ensure that the consultation process remains effective and responsive even as the number of nodes increases. Exploring solutions such as distributed computing, optimized query strategies, and hierarchical consultation frameworks could be potential pathways to making this paradigm scalable and efficient for real-world applications.
[…]
Appendix A Example Prompts
We display 2 example node prompts on datasets Cora and Wisconsin respectively in Table 7 for reproducibility.
Table 7. Example Prompts for Cora and Wisconsin.
Datasets Prompts
Cora Given a citation graph: { “id”:“cora_356”, “graph”: { “node_idx”: 356, “node_list”: [356, 190, 510, 519, 2223, 192], “one_hop_neighbors”: [190, 510], “two_hops_neighbors”: [192, 519, 2223], “node_label”: [“Neural_Networks”, “Neural_Networks”, “Probabilistic_Methods”, “Neural_Networks”, “Neural_Networks”, “Neural_Networks”] }}, where the 0th node is the target paper, with the following information: Title: [Title Text], Abstract: [Abstract Text]. And in the ‘node label’ list / ‘GNN-predicted node label’ list, you’ll find the human-annotated / GNN-predicted subcategories corresponding to the neighbors within two hops of the target paper as per the ‘node_list’.
Wisconsin Given a webpage link graph: { “id”: “wisconsin_62”, “graph”: { “node_idx”: 62, “node_list”: [62, 166, 189, 165, 84], “one_hop_neighbors”: [166, 189], “two_hops_neighbors”: [84, 165], “node_label”: [“course”, “student”, “student”, “student”, “faculty”] }}, where the 0th node is the target webpage, with the following content: [webpage content text]. And in the ‘node label’ list / ‘GNN-predicted node label’ list, you’ll find the human-annotated / GNN-predicted subcategories corresponding to the neighbors within two hops of the target paper as per the ‘node_list’.
Question: Which category does this webpage belong to? Give the most likely category of this webpage directly, choosing from “course”, “faculty”, “student”, “project”, “staff”. Ensure that your response can be parsed by Python json, using the following format as an example: {“classification result”: “student”, “explanation”: “your explanation for your classification here”,}. Ensure that the classification result must match one of the given choices.
Appendix B more baselines of different LLMs-for-Graphs Paradigms
To further show the effectiveness of the LLMs-as-Consultants paradigm, we conducted more experiments on Arxiv (Lim et al., 2021) dataset with more baselines of both LLMs-as-Predictors and LLMs-as-Enhancers paradigms as Table 8 shows. Note that Arxiv is a semi-homophilic graph with 169,343 nodes and 31,166,24 edges, among which 37% nodes are heterophilic.
GraphGPT (Tang et al., 2023) and InstructGLM (Ye et al., 2023) are LLMs-as-Predictors methods. The results of GraphGPT and InstructGLM on Arxiv dataset are directly borrowed from their papers.
SimTeG (Duan et al., 2023) and TAPE (He et al., 2023) are both LLMs-as-Enhancers methods. SimTeG involved LLM fine-tuning and TAPE prompted LLM for predictions and explanations to enhance the node features. For SimTeG, we trained GNNs with the released embeddings of Arxiv. For TAPE, we reproduced the prompting with Vicuna-7b-v1.5 to make the comparison fair. Note that SimTeG finetuned LLMs on all the training nodes and TAPE enhanced every node embedding, while our method only consults LLMs with some uncertain nodes, requiring less time and memory, but still performs on par with the baselines.
Table 8. Performance comparison with more baselines of LLMs-as-Predictors and LLMs-as-Enhancers paradigm on Arxiv.
Paradigm Method Acc
LLMs-as-Predictors GraphGPT 0.7511
LLMs-as-Predictors InstructGLM 0.7570 ± 0.0012
LLMs-as-Enhancers MLP+SimTeG 0.7432 ± 0.0023
LLMs-as-Enhancers MLP+TAPE 0.7461 ± 0.0029
LLMs-as-Consultants MLP+LOGIN 0.7476 ± 0.0036
LLMs-as-Enhancers GraphSAGE+SimTeG 0.7624 ± 0.0050
LLMs-as-Enhancers GraphSAGE+TAPE 0.7558± 0.0029
LLMs-as-Consultants GraphSAGE+LOGIN 0.7609± 0.0041